Session: Nov 2003 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (orfei at Fri Oct 24 11:04:35 2003) Success :

Torun (K.Borkowski at Fri Oct 24 11:36:12 2003) Failure : RECO
A problem with servomechanism of the reel driving motors in our P&G recorder.

Jodrell_Bank (Alastair Gunn at Fri Oct 24 12:58:08 2003) Minor failures :
Problems with procedures means that probably only pass 1 is good. Pass 2 maybe okay but passes 3 and 4 probably overwrote pass 2. Try pass 1 first. Problem now fixed.

Westerbork (tony foley at Fri Oct 24 14:07:03 2003) Minor failures : PHAS
We had big problems starting the backend. Some were due to static electricity in this dry weather. We have since laid an anti-static mat near the backend Data before 10:20UT must be bad. mark5 scans were send to disk ready for the real-time correlator

Effelsberg (A. Kraus at Sat Oct 25 14:54:28 2003) Minor failures : TSYS
In the first six minutes the IF-levels were not ok.

Noto (P. Cassaro at Sun Oct 26 08:06:34 2003) Success :